What are the odds of coin flip having the same results 8 times in a row?
1 in 27, or 1 in 128, or 0.0078125.
Note: the stated equation is 1 in 27, not 1 in 28, because the question said "same results 8 times in a row". If the question said "heads, 8 times in a row", then the answer would be 1 in 28, but since it said "same results", you have to consider that the second through the eighth flip must match the first flip, and that is 1 in 27.
What is the sample space for rolling a pair of dice?
The sample space is the following set:
{(1. 1), (1, 2), (1, 3), (1, 4), (1, 5), (1, 6),
(2. 1), (2, 2), (2, 3), (2, 4), (2, 5), (2, 6),
(3. 1), (3, 2), (3, 3), (3, 4), (3, 5), (3, 6),
(4. 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6),
(5. 1), (5, 2), (5, 3), (5, 4), (5, 5), (5, 6),
(6. 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6)}
What is the sample space for picking a random card from a deck?
The sample space is each card in the deck.
What does contextual status mean?
status that aren't mutually exclusive. it refers to person who involve in multiple status. example: some women can be both mother and a businesswoman
What are the possible outcomes for the throw of 2 dice?
When two dies are thrown, there are 36 possible outcomes. The first number represents the outcome of the first die and the second number represents the outcome of the second die.
(1,1),(1,2),(1,3),(1,4),(1,5),(1,6)
(2,1),(2,2),(2,3),(2,4),(2,5),(2,6)
(3,1),(3,2),(3,3),(3,4),(3,5),(3,6)
(4,1),(4,2),(4,3),(4,4),(4,5),(4,6)
(5,1),(5,2),(5,3),(5,4),(5,5),(5,6)
(6,1),(6,2),(6,3),(6,4),(6,5),(6,6)
What is the probability that A date picked at random is February 30?
The probability that a date picked at random is February 30th is zero, as there is no February 30th, even in leap years.
Six times the number of different outcomes on the spinner.
How many different ways can five books be arranged on a shelf?
The number of permutations of five things taken five at a time is five factorial, or 120.
What does large mean absolute deviation means?
Large means something or an object that is big in size.
How are experimental probability and theoretical experiment alike?
They are both estimates of the probability of outcomes that are of interest. Experimental probabilities are derived by repeating the experiment a large number of times to arrive at these estimates whereas theoretical probabilities are estimates based on a mathematical model based on some assumptions.
When two dice are rolled what is the probability that both will be a five?
if you multiply the number of rolls you did you will get the answer
z = (75 - 85)/5 = -10/5 = -2
What is mean by cardinality of sample space?
no of possibilities
for example tossind a fair coin then
the cardinality of sample space is 2
What is the mean mode median standard deviation and standard error of 10 24 35 44 10 and 35?
Mean: 26.33
Median: 29.5
Mode: 10, 35
Standard Deviation: 14.1515
Standard Error: 5.7773
It is 0.0606
The standard z-score such that 80 percent of the distribution is below to the left of this value is?
z = 0.8416
What is the probability of toss two heads with 2 pennies?
There are four different ways that two coins can land:
T T
T H
H T
H H.
Only one of them is two heads, so if the coins are fair, then
the probability is 1/4 = 25% .
What are the similarities of a theoretical and empirical probabilities?
They are both measures of the likelihood of specified events.
Why is t score equal to z score in a normal distribution?
Because as the sample size increases the Student's t-distribution approaches the standard normal.
What is the probability of drawing an ace or a king?
There are 52 cards in a deck
there are 4 aces and 4 kings which makes a total of 8 kings and aces.
Assuming that the deck is full and shuffled the probability of drawing an aces or a king is 8/52 which simplifies to 2/13
How many different ways can you arrange 20 photos in an array?
Assuming that the arrays need to be rectangular, the answer is 14,597,412,049,059,800,000 or approx 15 quintillion.